# Renormalization group improved cosmology in the presence of a stiff matter era

### v1

@@ III Modified equation of state using the relativistic thermodynamical approach
where $s$ denotes the entropy density and $\varepsilon$ denotes the energy density for a system. If $N$ is the number of particles inside a box of volume $V$ with each particles having a mass $m$, then dividing the left hand side and right hand side of the above equation with $Nm$, we obtain the local form of the first law of thermodynamics [37, 34]
@@
Making use of the above equation in eq.(7), we obtain the following relation [34]
@@
with $\mathcal{A}$ denoting the integration constant. Instead of fixing the constant using the current time values of $\rho$ and $a$, we focus mainly at the fixed point time as we are mainly focussed in investigating the dynamics of the universe before the fixed point regime or the Planck time ($t_{\text{Pl}}$). In the current analysis $t_{*}^{UV}=t_{\text{Pl}}$. At $t=t_{*}^{UV}$, $\rho=\rho_{*}^{UV}$ and $a=a_{*}^{UV}$ [^3] For removing cluttering in the expressions, we write denote the scale factor and the energy density of the universe at the fixed time as $a_{*}^{UV}=a_{*}$ and $\rho_{*}^{UV}=\rho_{*}$.. Hence, we find that the integration constant has the value $\mathcal{A}={a_{*}^{UV}}^{3}\rho_{*}^{UV}=a_{*}^{3}\rho_{*}$. We can now recast eq.(25) as
@@
Within the fixed time regime (considering the universe is always expanding) ($t\ll t_{\text{Pl}}$), $a\ll a_{*}$ but for our calculation we shall make the range of the scale factor $a(t)$ a bit flexible by considering eq.(s)(9,10) to hold when $a<a_{*}$ as well.
Again from eq.(23), we know that $P=P(\rho)$ and we obtain the analytical form of the energy density as a function of pressure and rest-mass density of the fluid as [34]
@@
According to the model proposed by Zel’dovich [29, 30], the complete equation of state describing the cosmological evolution process reads
@@
and the energy density was expressed as $\varepsilon(\rho)=\rho+K_{2}\rho^{2}$ which is same as eq.(30). The above expression for the energy density can be further augmented to the form given by [^4] For reference see [34].
@@
which also implies that $\Omega_{\text{s},*}+\Omega_{\text{rad},*}+\Omega_{\text{m},*}=1$ as $\varepsilon=\varepsilon_{*}$ when $a=a_{*}$ (for $t=t_{\text{Pl}}$).
IV RG flow improved Friedmann equations with a stiff matter era before the fixed point
In this section we shall investigate the renormalization flow improved Friedmann equations when there is a stiff matter era before the Planck time and we shall also discuss the evolution of the scale factor after the Planck time. Substituting $G(k)$, $\Lambda(k)$, and $\varepsilon(k)$ from eq.(s)(9,10,15), in eq.(11), we obtain the Friedmann equation in terms of $k$ as
@@ IV RG flow improved Friedmann equations with a stiff matter era before the fixed point
where we have made use of the identification $\lambda_{*}\equiv\lambda_{*}^{UV}$.
@@ IV.1 Stiff matter and radiation era
We shall now consider the universe with an early stiff matter and radiation era. For such a model, the energy density of the universe varies with the scale factor as
@@
If we define a new variable $\xi\equiv\frac{a}{a_{*}}$, then the minimum value that $\xi$ can attain is zero and the upper limit will always be less than unity as inside the fixed point regime $a<a_{*}$. Our primary aim now is to obtain the time evolution of the scale factor $a(t)$. In terms of the new variable $\xi$, we can express $H$ as
@@
Now $\xi<1$, and as a result it is always quite straightforward to obtain a series expansion of the Gauss hypergeometric function in the above equation. The problem with the exact solution is that it is impossible to obtain $\xi$ in terms of the time $t$. However, upto leading order contribution, we obtain $a$ in terms of $t$ as
@@
We find out that in the fixed point regime the leading order contribution to the evolution of the scale factor will come from the stiff-matter fluid with a time dependece proportional to $t^{\frac{2}{3}}$. In order to obtain the contribution due to the radiation term, one needs to consider higher order contributions in $\xi$ in eq.(44). Expanding the right hand side of eq.(44) in powers of $\xi$, we can express eq.(44) as
@@
From Fig.(1), we observe that with increase in time $\frac{a_{+}(t)}{a_{*}}$ decreases which goes against the well-known behaviour of the universe leading to a phantom behaviour. If one plots the energy density from eq.(38) using $a_{+}(t)$, one observes that the energy density increases with time which indicates a phantom-like behaviour for this particular time dependence of the scale factor. Whereas we can see that both $a_{-}(t)$ and $a_{0}(t)$ shows quite a convincing behaviour as with increase in time the scale factor increases. It is though important to understand that as $\Omega_{\text{rad},*}$ approaches zero, $a_{-}(t)\rightarrow a_{0}(t)$, which is an expected behaviour. We therefore consider $a_{-}(t)$ to show the most convincing behaviour of the scale factor and set it as $a(t)$. In principle, $a_{0}(t)$ shows the time-dependence of the scale factor in absence of any radiation as can be obtained directly by setting $\Omega_{\text{rad},*}=0$ in eq.(44) (In such a scenario $\Omega_{\text{s},*}=1$). Using $a_{0}(t)$ and $a_{-}(t)$ from eq.(s)(45,48) in eq.(38), we get the time depence of the energy density of the universe. The $a_{0}(t)$ denotes the energy density of the universe in absence of radiation defined as $\varepsilon_{0}(t)$, whereas $a(t)$ gives the energy density in presence of both stiff matter and radiation component before the fixed time. It is although important to note that in Fig.(1), $\frac{a}{a_{*}}$ has exceeded unity. From the $\frac{a}{a_{*}}=1$ condition, we can obtain the value of $H_{*}t$, which in the current case is obtained to be $H_{*}t\simeq 0.68$. Hence, upto $\frac{a}{a_{*}}=1$, we can plot the scale factor versus the dimensionless times in Fig.(2).
@@
Now the above inequality only points toward the validity of the solution presented in eq.(48). Next, we shall investigate the case when there is a matter dominated era present with the radiation and stiff matter era in the very early universe.
@@ IV.2 Stiff matter, radiation, and dark matter
We consider that the universe consists of a stiff matter era, radiation era, and a dark matter era and the corresponding energy density is given by eq.(36). At the fixed point, $\varepsilon\rightarrow\varepsilon_{*}$, which leads to the condition
@@
with $\zeta(t)$ being function of time which needs to be determined. Keeping terms upto $\mathcal{O}\left(\frac{\Omega_{\text{m}*}}{\Omega_{\text{s},*}},\frac{\Omega_{\text{rad}*}}{\Omega_{\text{s},*}}\right)$, and substituting the above solution back in eq.(54), we obtain the analytical form of $\zeta(t)$ as
@@
We now plot the above result with respect to the case when there is no dark matter or radiation domination in the universe implying $\Omega_{\text{s},*}=1$. For the current case, we set, $\Omega_{\text{s},*}=0.7$, $\Omega_{\text{rad},*}=0.2$, and $\Omega_{\text{m},*}=0.1$. We now plot the dimensionless scale factor against the dimensionless time in Fig.(4).
@@
In the above plot, we observe that in the very early time stiff matter dominates the evolution of the scale factor. In the later part, radiation and dark matter domination becomes significant. In Fig.(4), the dotted red line denotes the $\frac{a(t)}{a_{*}}=1$ line. We now plot the dimensionless energy densities for the three cases when there is stiff-matter+radiation+dark-matter, stiff-matter+radiation, and stiff-matter only in the universe against the dimensionless time. For the stiff-matter+radiation+dark-matter case, the energy density is described by $\varepsilon_{\text{rm}}(t)$, for the stiff-matter case the energy density of the universe is described by $\varepsilon_{\text{r}}(t)$, and for the stiff-matter only case, the energy density is described by $\varepsilon_{0}(t)$ in Fig.(5). In Fig.(5), for the $\varepsilon_{\text{rm}}$ case $\Omega_{\text{s},*}=0.6$, $\Omega_{\text{rad},*=0.25}$, and $\Omega_{\text{m},*}=0.15$, whereas for the $\varepsilon_{\text{r}}$ case, $\Omega_{\text{s},*}=0.6$ and $\Omega_{\text{rad},*}=0.4$.

### v3

@@ III Modified equation of state using the relativistic thermodynamical approach
where $s$ denotes the entropy density and $\varepsilon$ denotes the energy density for a system. If $N$ is the number of particles inside a box of volume $V$ with each particles having a mass $m$, then dividing the left hand side and right hand side of the above equation with $Nm$, we obtain the local form of the first law of thermodynamics Weinberg ; Chavanis0
@@
Making use of the above equation in the standard continuity equation in FLRW cosmology, we obtain the following relation Chavanis0
@@
with $\mathcal{A}$ denoting the integration constant. Instead of fixing the constant using the current time values of $\rho$ and $a$, we focus mainly about the Planck time regime as we are mainly focused in investigating the dynamics of the universe before the Planck time ($t_{\text{Pl}}$). In the current analysis $t_{*}^{UV}=t_{\text{Pl}}$. At $t=t_{*}^{UV}$, $\rho=\rho_{*}^{UV}$ and $a=a_{*}^{UV}$ [^4] For removing cluttering in the expressions, we write denote the scale factor and the energy density of the universe at the fixed time as $a_{*}^{UV}=a_{*}$ and $\rho_{*}^{UV}=\rho_{*}$.. Hence, we find that the integration constant has the value $\mathcal{A}={a_{*}^{UV}}^{3}\rho_{*}^{UV}=a_{*}^{3}\rho_{*}$. We can now recast eq.(20) as
@@
Within the fixed time regime (considering the universe is always expanding) ($t\ll t_{\text{Pl}}$), $a\ll a_{*}$ but for our calculation we shall make the range of the scale factor $a(t)$ a bit flexible by considering eq.(s)(3,4) to hold when $a<a_{*}$ as well. Eq.(21) is different from the usual structure as the reference point is the fixed-point regime instead of the current-time as is usually done in the literature Weinberg ; WeinbergCosmo ; Ryden .
Again from eq.(18), we know that $P=P(\rho)$ and we obtain the analytical form of the energy density as a function of pressure and rest-mass density of the fluid as Chavanis0
@@
According to the model proposed by Zel’dovich Zeldovich ; Zeldovich2 , the complete equation of state describing the cosmological evolution process reads
@@
and the energy density was expressed as $\varepsilon(\rho)=\rho+K_{2}\rho^{2}$ which same as eq.(25). The above expression for the energy density can be further simplified to the form given by [^5] For reference see Chavanis0 .
@@
which also implies that $\Omega_{\text{s},*}+\Omega_{\text{rad},*}+\Omega_{\text{m},*}=1$ as $\varepsilon=\varepsilon_{*}$ when $a=a_{*}$ (for $t=t_{\text{Pl}}$). The above energy density although have a structure similar to the one discussed in Chavanis0 , it uses Planck time ($t_{*}\equiv t_{Pl}$) as the time of reference whereas for the case in Chavanis0 , the reference time is the current time of our universe ($t_{0}$) [^6] We have discussed the analytical form of the energy density (used in this analysis) in details to truly highlight the time of reference used here in contrast to the earlier analyses Chavanis0 ; Chavanis1 ; Chavanis2 ..
IV RG flow improved Friedmann equations with a stiff matter era above the Planck scale ($k>m_{Pl}$)
In this section we shall investigate the renormalization flow improved Friedmann equations when there is a stiff matter era before the Planck time and we shall also discuss the evolution of the scale factor after the Planck time. Substituting $G(k)$, $\Lambda(k)$, and $\varepsilon(k)$ from eq.(s)(3,4,10), in eq.(5), we obtain the Friedmann equation in terms of $k$ as
@@ IV RG flow improved Friedmann equations with a stiff matter era above the Planck scale ($k>m_{Pl}$)
where we have made use of the identification $\lambda_{*}\equiv\lambda_{*}^{UV}$. From eq.(32), we can see that $H=\frac{\dot{a}}{a}$ is dependent upon $k$. This dependence upon $k$ originates due to the flow of the cosmological constant $\Lambda$. Therefore, $H$ has no constant contribution.
@@ IV.1 Stiff matter and radiation era
Just after the inflationary phase, the universe enters into the reheating epoch. This reheating era heats up the cold and inflated universe. It is therefore largely considered that the radiation era existed just after the inflation era near to the beginning of nucleosynthesis. The inflation era is assumed to happen between $10^{-33}$ sec and $10^{-32}$ sec which is way after the Planck time of the universe. It was assumed in Zeldovich ; Zeldovich2 , that the universe consisted of a cold baryon gas which obeys a stiff equation of state and contributed to the evolution of the universe in the post-inflation period and just before the radiation era. In OdintsovOikonomou , a stiff matter era was considered in the post-inflationary period in the case of $f(R)$ gravity. In case of going with the traditional way of incorporating stiff-matter era, we assume the presence of a dark-fluid which behaves as a stiff-matter fluid inside the scaling regime when $t<t_{Pl}$. This allows one for a very early “power-law” expansion of the universe. In the cosmological renormalization group improved models, before the Planck time, the energy density is very high and as a result it is quite clearly understandable that any fluid existing at such an early time will have pressure so high that the speed of sound will approach the speed of light. As argued in Chavanis0 , the universe at earlier times is considered to be filled with a dark-fluid. In the $0<t\leq t_{Pl}$, we consider the energy density of the dark-fluid to have a $a^{-6}$ dependence and a $a^{-4}$ dependence. We incorporate a radiation-like part in the energy density which is expected to dominate after the Plack time. In contrast to the general cosmological models here, one considers a single fluid which shows different behaviour with respect to the change in time. We shall now therefore consider the universe with an early stiff matter and radiation era. For such a model, the energy density of the universe varies with the scale factor as
@@
If we define a new variable $\xi\equiv\frac{a}{a_{*}}$, then the minimum value that $\xi$ can attain is zero and the upper limit will always be less than unity as inside the fixed point scaling regime $a<a_{*}$. Our primary aim now is to obtain the time evolution of the scale factor $a(t)$. In terms of the new variable $\xi$, we can express $H$ as
@@
In the $t<t_{*}(t_{Pl})$ regime, $\xi<1$, and as a result it is always quite straightforward to obtain a series expansion of the Gauss hypergeometric function in the above equation considering $\xi$ to be the small parameter. The problem with the exact solution is that it is impossible to obtain $\xi$ in terms of the time $t$. However, upto leading order contribution, we obtain $a$ in terms of $t$ as
@@
We find out that in the fixed point scaling regime the leading order contribution to the evolution of the scale factor will come from the stiff-matter fluid with a time dependece proportional to $t^{\frac{2}{3}}$. In order to obtain the contribution due to the radiation term, one needs to consider higher order contributions in $\xi$ in eq.(39). Expanding the right hand side of eq.(39) in powers of $\xi$, we can express eq.(39) as
@@
From Fig.(1), we observe that with increase in time $\frac{a_{+}(t)}{a_{*}}$ decreases which goes against the well-known behaviour of the universe leading to a phantom behaviour. If one plots the energy density from eq.(33) using $a_{+}(t)$, one observes that the energy density increases with time which confirms this phantom-like behaviour for this particular time dependence of the scale factor. Whereas we can see that both $a_{-}(t)$ and $a_{0}(t)$ shows quite a convincing behaviour as with increase in time the scale factor increases. It is though important to understand that as $\Omega_{\text{rad},*}$ approaches zero, $a_{-}(t)\rightarrow a_{0}(t)$, which is an expected behaviour. We therefore consider $a_{-}(t)$ to show the most convincing behaviour of the scale factor and set it as $a(t)$. In principle, $a_{0}(t)$ shows the time-dependence of the scale factor in absence of any radiation as can be obtained directly by setting $\Omega_{\text{rad},*}=0$ in eq.(39) (In such a scenario $\Omega_{\text{s},*}=1$). Using $a_{0}(t)$ and $a_{-}(t)$ from eq.(s)(40,43) in eq.(33), we get the time depence of the energy density of the universe. The $a_{0}(t)$ denotes the energy density of the universe in absence of radiation defined as $\varepsilon_{0}(t)$, whereas $a(t)$ gives the energy density in presence of both stiff matter and radiation component before the fixed time. It is although important to note that in Fig.(1), $\frac{a}{a_{*}}$ has exceeded unity. From the $\frac{a}{a_{*}}=1$ condition, we can obtain the value of $H_{*}t$, which in the current case is obtained to be $H_{*}t\simeq 0.68$. Hence, upto $\frac{a}{a_{*}}=1$, we can plot the scale factor versus the dimensionless times in Fig.(2).
@@
Now the above inequality only points toward the validity of the solution presented in eq.(43). Some significant changes of this flow improved cosmology with respect to the standard FLRW cosmology are in order.
  1. 1.
    The dependence of the scale factor in the stiff-matter dominated era goes as $t^{\frac{2}{3}}$ for the renormalization group improved cosmological model whereas for the standard FLRW cosmology the scale factor has the time dependence given by $t^{\frac{1}{3}}$ Chavanis0 . This implies that for our case the expansion in the stiff-matter dominated era will be enhanced during the fixed point scaling regime which is the regime between the Big Bang and the Planck time.
  2. 2.
    Again, in the dimensionless time parameter $H_{*}t\equiv H(t_{Pl})t$, the Hubble’s constant will be much higher as $H(t_{Pl})\gg H_{0}$ leading to a more enhanched “power-law” expansion compared to the standard FLRW case inside the fixed point scaling regime before Planck time.
  3. 3.
    The flow improved cosmology portrays clearly that the domination of the stiff matter era ended when the relative scale factor becomes unity at the Planck time whereas for the FLRW cosmology the clear transition point between stiff-matter dominated era and radiation era is unclear Chavanis0 .
In the next subsection, we shall also incorporate a dark-matter like part in the energy density and observe whether it has any substantial effect inside the fixed point scaling regime.
@@ IV.2 Stiff matter, radiation, and dark matter
We consider that the universe was filled with a dark fluid which has a stiff matter-like, radiation-like, and dark matter-like contributions and the corresponding energy density is given by eq.(31). At the fixed point we get $\varepsilon\rightarrow\varepsilon_{*}$, which leads to the condition
@@
One important thing to observe from the above equation is that the right hand side has no time independent constant terms and as a result, in this flow-improved cosmology, an early-time inflation is not possible. This conclusion is in line with the arguments presented in BonanoReuter2 . Inflation can be realized by introducing new equations of state which we shall present in section (VII).
@@
with $\zeta(t)$ being function of time which needs to be determined. Keeping terms upto $\mathcal{O}\left(\frac{\Omega_{\text{m}*}}{\Omega_{\text{s},*}}\right)$ and $\mathcal{O}\left(\frac{\Omega_{\text{rad}*}}{\Omega_{\text{s},*}}\right)$, and substituting the above solution back in eq.(49), we obtain the analytical form of $\zeta(t)$ as
@@
We now plot the above result with respect to the case when there is no dark matter or radiation-like component of the dark fluid in the universe implying $\Omega_{\text{s},*}=1$. For the current case, we set, $\Omega_{\text{s},*}=0.7$, $\Omega_{\text{rad},*}=0.2$, and $\Omega_{\text{m},*}=0.1$. It is to be noted that the values that we have chosen here are in sharp contrast to the those that have been taken in Chavanis0 . In particular, we have considered $\Omega_{s,*}>\Omega_{\text{rad},*}>\Omega_{\text{m},*}$. The physical motivation behind this consideration lie in the fact that near the fixed point regime the dark-fluid primarily behaves as a stiff-matter whereas radiation and matter like components remain primarily latent. We now plot the dimensionless scale factor against the dimensionless time in Fig.(4).
@@
In the above plot, we observe that in the very early time stiff matter-like component dominates the evolution of the scale factor. In the later part, there is slight contribution due to radiation-like and dark matter-like component of the dark fluid. As expected, the dark-matter or radiation like component has almost no significant effect inside the fixed point regime implying that for $t<t_{*}$, the dark-fluid behaves primarily as a stiff matter, and dominates the expansion of the universe. Hence, without loss of any generality, one can consider eq.(31) as the most general energy density for a dark-fluid in a renormalization flow improved cosmology. The rationale behind taking the other components, namely, the radiation and dark-matter is that these components will start to play an important role at later epochs as will be discussed subsequently. In Fig.(4), the dotted red line denotes the $\frac{a(t)}{a_{*}}=1$ line. We now plot the dimensionless energy densities for the three cases when there is stiff-matter+radiation+dark-matter, stiff-matter+radiation, and stiff-matter only in the universe against the dimensionless time. For the stiff-matter+radiation+dark-matter case, the energy density is described by $\varepsilon_{\text{rm}}(t)$, for the stiff-matter case the energy density of the universe is described by $\varepsilon_{\text{r}}(t)$, and for the stiff-matter only case, the energy density is described by $\varepsilon_{0}(t)$ in Fig.(5). In Fig.(5), for the $\varepsilon_{\text{rm}}$ case $\Omega_{\text{s},*}=0.6$, $\Omega_{\text{rad},*}=0.25$, and $\Omega_{\text{m},*}=0.15$, whereas for the $\varepsilon_{\text{r}}$ case, $\Omega_{\text{s},*}=0.6$ and $\Omega_{\text{rad},*}=0.4$.